The volume V of a gas varies inversely as the pressure P in it. The volume of a gas in 240cm^3 under a pressure of 23kg/cm^2. What will be its volume under a pressre of 30kg/cm^2?: The volume V of a gas varies inversely as the pressure P in it. The volume of a gas in 240cm^3 under a pressure of 23kg/cm^2. What will be its volume under a pressre of 30kg/cm^2?
Answer by stanbon(18743)

The volume V of a gas varies inversely as the pressure P in it.
V = k/P
-------------------
The volume of a gas in 240cm^3 under a pressure of 23kg/cm^2.
240 = k/23
k = 5520
EQUATION for your problem:
V = 5520/P
---------------
What will be its volume under a pressre of 30kg/cm^2?
V = 5520/30
V = 184 cm^3
=================
Cheers,
Stan H.